(25) Tickets for a concert were sold to adults for $3 and to students for $2. If the total receipts were $824 and twice as many adult tickets as student tickets were sold, then how many of each were sold?


Found 2 solutions by monika_p, stanbon:
Answer by monika_p(71)

if x is the number of tickets sold for students for $2, then 2x is number of tickets sold for adults for $3. Now we can write equation:
2*x+3*2x=824
8x=824
x=103 number of tickets sold for students, then for adults was sold 206 tickets


Answer by stanbon(48542)

Tickets for a concert were sold to adults for $3 and to students for $2. If the total receipts were $824 and twice as many adult tickets as student tickets were sold, then how many of each were sold?
-------------------------------------------------
Equations:
Value Equation: 3a + 2s = 824
Quantity Eq...: a - 2s = 0
------------------------------
Add the two equations to get:
4a= 824
a = 206 (# of adult tickets sold)
s = 103 (# of student tickets sold)
---------------
Cheers,
Stan H.